# LongVILA: Scaling Long-Context Visual Language Models for Long Videos

### v3

@@ Abstract
Long-context capability is critical for multi-modal foundation models, especially for long video understanding. We introduce LongVILA, a full-stack solution for long-context visual-language models by co-designing the algorithm and system. For model training, we upgrade existing VLMs to support long video understanding by incorporating two additional stages, i.e., long context extension and long supervised fine-tuning. However, training on long video is computationally and memory intensive. We introduce the long-context Multi-Modal Sequence Parallelism (MM-SP) system that efficiently parallelizes long video training and inference, enabling 2M context length training on 256 GPUs without any gradient checkpointing. LongVILA efficiently extends the number of video frames of VILA from 8 to 1024, improving the long video captioning score from 2.00 to 3.26 (out of 5), achieving 99.5% accuracy in 1400-frame (274k context length) video needle-in-a-haystack. LongVILA-8B demonstrates consistent accuracy improvements on long videos in the VideoMME benchmark as the number of frames increases. Besides, MM-SP is 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than ring sequence parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron with context parallelism + tensor parallelism. Moreover, it seamlessly integrates with Hugging Face Transformers.
@@ 1 Introduction
While some works have enabled long-context Vision-Language Models (VLMs) (Lin et al., 2023b; Weng et al., 2024), they employ simplified approaches rather than offering a comprehensive solution. For instance, LongVA (Zhang et al., 2024a) depends on long-context LLMs and trains models on short-context data. LongVLM (Weng et al., 2024) utilizes token compression to circumvent context extension. These approaches sidestep more challenging issues, such as the development of a robust long-context multi-modal training framework and corresponding dataset design.
@@
A full-stack design is crucial for long-context Vision-Language Models (VLMs). Training large models is typically a complex, systematic endeavor that demands both data engineering (Betker et al., 2023; Ouyang et al., 2022; Zhou et al., 2024) and system-software co-design (Lepikhin et al., 2020; Chowdhery et al., 2023; Shoeybi et al., 2019; Brown et al., 2020; Dehghani et al., 2023). Unlike text-only LLMs, VLMs (e.g., LLaVA (Liu et al., 2023b)) often require distinct model architectures and flexible distributed training strategies. Additionally, long-context modeling necessitates not only long-context data to fully utilize the model’s capabilities (Fu et al., 2024b; Chen et al., 2023) but also infrastructure capable of supporting memory-intensive long-context training (Li et al., 2021; Jacobs et al., 2023; Li et al., 2023a). Therefore, a full-stack design, encompassing training pipeline and system, is indispensable for long-context VLMs.
@@
Figure 1: The LongVILA training pipeline for long video understanding. In Stages 1 through 3, the process starts with alignment, pre-training, and supervised fine-tuning. In Stage 4, the model undergoes mid-training context extension. Finally, in Stage 5, the model is fine-tuned for long video understanding with Multi-Modal Sequence Parallelism (MM-SP). While this example uses VILA, the process is adaptable to other VLMs. We train LongVILA models to support 1024 video frames in Stage 5.
@@
In this work, we introduce LongVILA, a comprehensive solution for long-context VLMs. For training pipeline, we implement a five-stage training curriculum as Figure 1: (1) multi-modal alignment, (2) large-scale pre-training, (3) short supervised fine-tuning, (4) context extension for LLMs, and (5) long supervised fine-tuning. For training system, we establish an efficient and user-friendly framework, namely Multi-Modal Sequence Parallelism (MM-SP), which supports training memory-intensive long-context VLMs. For inference, MM-SP addresses the challenge of KV cache memory usage, which becomes a bottleneck with very long sequences, by deploying an inference mode within our MM-SP framework to support long-context multi-modal language deployment.
@@
By efficiently increasing the number of video frames using LongVILA, the performance on VideoMME (Fu et al., 2024a) and long video captioning tasks consistently improves (Figure 2). The LongVILA model, trained on 1024 frames, achieves 99.5% accuracy in the needle-in-a-haystack experiments with 1400 frames, equivalent to a context length of 274k tokens. Our MM-SP system can efficiently scale the context length up to 2 million tokens without gradient checkpointing, achieving 2.1$\times$ to 5.7$\times$ speedup compared to ring sequence parallelism, and 1.1$\times$ to 1.4$\times$ compared to Megatron with context parallelism + tensor parallelism.
@@ 2 Background and Related work
There are two predominant designs for VLMs: the encoder-decoder architecture (e.g.,, LLaVA (Liu et al., 2023b), PaLM-E (Driess et al., 2023)) and the decoder-only architecture (e.g.,, Fuyu (Bavishi et al., 2023), Chameleon (Team, 2024)). Encoder-Decoder VLMs connect the vision encoder to the LLM decoder through a multi-modal projector. Certain multi-modal projectors, such as spatial pooling and Q-former, significantly reduce the number of tokens per image or video frame, thereby lowering the computational burden on the LLM decoder. In contrast, decoder-only LLMs typically process raw patches as input without hierarchical token pooling, making it more challenging to reduce the token count for each image or frame. In this work, we build on VILA (Lin et al., 2023b) as our foundation. It is worth noting that enhanced variants of VILA exist, such as VILA2(Fang et al., 2024) for improved performance and X-VILA(Ye et al., 2024) for cross-modality understanding, reasoning, and generation. For our model architecture and training pipeline, we adhere to the standard VILA-1.5 version.
@@ 3 LongVILA Training Pipeline
As shown in Figure 1, in our pipeline, there are five training stages, i.e., Stage 1: multi-modal alignment, Stage 2: large-scale pre-training, Stage 3: supervised fine-tuning, Stage 4: context extension for LLM, Stage 5: long supervised fine-tuning. Stage 1, 2, and 3 are following VILA Lin et al. (2023b), to firstly bridge the gap between LLM and vision encoder, and then pre-training on larger datasets. In Stage 1, only the multi-modal projector is trainable with others frozen. In Stage 2, we frozen the vision encoder and training LLM and the multi-modal projector. In Stage 3, we fully fine-tune the model for short data instruction following, e.g., image and short video datasets. Afterwards, we extend the context length of LLM with text-only dataset in a continued pre-training manner in Stage 4. In Stage 5, we adopt our MM-SP system (§4) to enhance the instruction following abilities by long video supervised fine-tuning. It is noted that all parameters are trainable in the final stage.
@@ 3.1 Stage1&2&3: Alignment, Pre-training, and Short Supervised Fine-tuning
Before training at scale, we first use open-sourced image and video caption datasets to train the multi-modal projector in stage (1) to conduct the multi-modal alignment. Note that, following (Lin et al., 2023b), both vision encoder and LLM decoder are frozen at this stage. After that, we conduct large-scale pre-training to learn general multi-modal capacity at scale. To improve the quality of large open-sourced datasets, we follow VILA2 (Fang et al., 2024) to relabel COYO-25M (Lin et al., 2023b; Byeon et al., 2022) with VILA-1.5-40B (Lin et al., 2023b). The supervised fine-tuning process incorporates mixed data types, including both images and videos. For short video comprehension, we utilize open-source video instruction-following datasets, e.g., YouCook2 Zhou et al. (2018) and ShareGPTVideo Zhang et al. (2024b).
@@ 3.2 Stage4: Context Extension for LLMs
Our empirical research indicates that extending the context length of LLMs is essential prior to engaging in supervised fine-tuning with long video datasets. Following Stage 2 of our methodology, we execute a continuation of pre-training on the LLM to enhance its context length to 262,144, utilizing a total of 17B tokens. We employ a progressive training schedule, incrementally increasing the context length from 8,192 to 65,536, and ultimately to 262,144, utilizing the SlimPajama dataset (Soboleva et al., 2023) in accordance with the methodology outlined by (Fu et al., 2024c).
@@
Figure 4: The proportion of question and video categories in our LongVILA__\__sft dataset. We have 15,292 videos in total. For each video, there are one sample for captioning and the other question. The left pie chart shows the proportion of video categories. The right pie chart shows the proportion of question categories.
@@ 3.3 Stage5: Long Supervised Fine-tuning
Long video instruction following To facilitate the fine-tuning of long videos, we also constructed a new, dedicated dataset for long video training, consisting of 15,292 videos for each one. We use the original long videos from the Shot2Story dataset (Han et al., 2023). Each video includes different questions and answers: one for generating captions and another for answering questions, enabling diverse applications in video understanding. Figure 3 illustrates the process for generating instruction-following datasets from long videos. Initially, the long video is segmented into shorter clips, each approximately 10 seconds in duration. These clips are then independently annotated with descriptive captions utilizing the VILA-1.5 model. Subsequently, an LLM is employed to generate question-and-answer pairs derived from the captions of these clips. The generated questions encompass summarization and other queries relevant to the comprehensive understanding of long video content.
@@
Once we acquired the long video dataset, applying it for supervised fine-tuning introduced new challenges, primarily due to the substantial number of frames in each sample—often ranging in the hundreds or even thousands. For instance, a single sequence from 1400 video frames can encompass around 274k tokens. Existing data-parallel training systems struggle to handles such extensive contexts. To address this limitation, we developed the MM-SP system (§4) to efficiently train long-context visual language models.
@@ 4 Multi-Modal Sequence Parallelism
Figure 5: Sharding strategy and communication pattern of MM-SP. For sharding strategy, Ring SP is designed for text-only modalities, without optimization for the workload of an image encoder. Our MM-SP implements a novel sharding strategy that balances the computational load between the image encoder and the language modeling stages. For communication pattern, Ring SP (Liu et al., 2023a; Li et al., 2023a) (top) relies on P2P communication for both intra-node and inter-node settings, resulting in underutilization of intra-node bandwidth. MM-SP (bottom) adopts 2D-Attention (Fang & Zhao, 2024; Gu et al., 2024) mechenism which utilizes intra-node All-to-All (All2All) and inter-node Point-to-Point (P2P) communication to transfer keys and values (KV), enhancing the efficiency of intra-node NVLink utilization. The bandwidth is for H100.
4.1 Limitations of Existing Systems
@@ 4.2 Multi-Modal Sequence Parallelism Training Mode
Figure 6: Workflow of Multi-Modal Sequence Parallelism with a Batch Size (bs) of 4 and a Sequence Parallel Size (SP_Size) of 3. To facilitate sequence parallelism for multi-modal inputs, we developed a customized sharding strategy that ensures balanced workload distribution and compatibility with SP communucation.
@@ 4.3 Multi-Modal Sequence Parallelism Inference Mode
The model we developed through sequence parallelism training is capable of handling long-context multi-modal downstream tasks. However, the most commonly used inference system, built on HuggingFace Transformers, typically operates on a single GPU. This lack of distributed implementation limits the maximum sequence length that can be processed during inference. The most straightforward and user-friendly solution within HuggingFace Transformers is to use its pipeline parallelism inference feature, which shards a single model across multiple devices on a layer-by-layer basis (Huang et al., 2019; Narayanan et al., 2019). However, this approach is inefficient, as it only activates one GPU at a time. Additionally, it struggles to support long sequence lengths because the first device must store large tensors, such as input embeddings and images, creating a bottleneck that hampers the entire system’s performance.
To address these limitations, we implemented sequence parallelism for distributed inference in VLMs. Unlike the training mode, the inference system must additionally manage tensors, such as input tokens and position encodings, that progressively change during the decoding phase (Yu et al., 2022). The system also needs to detect signals from the machine holding the last token to terminate the distributed process appropriately. Compared to HuggingFace’s native pipeline parallelism strategy, our inference mode is more efficient, as all devices participate in computation simultaneously, accelerating the process by a factor proportional to the number of machines (Figure 7). Furthermore, it is scalable, with memory evenly distributed across devices, enabling support for longer sequences with additional machines.
@@ 5.1 Training and inference system
Table 2: Training system throughput comparison across various Megatron-LM strategies on 32 H100 GPUs, measured in time per iteration (seconds). ”OOM” indicates ”Out of Memory”
Sequence length Megatron-LM Ours
CP CP=4+TP=8 ZigZag-RingAttn Ulysses 2D Attention
320 K OOM OOM 23.57 10.70 11.12
288 K OOM OOM 20.24 8.68 8.65
256 K OOM OOM 17.54 6.98 7.04
224 K OOM OOM 15.22 5.47 5.53
192 K OOM OOM 12.97 4.15 4.24
160 K OOM OOM 10.83 3.02 3.11
128 K OOM 2.96 8.38 2.07 2.17
96 K 4.32 1.77 6.35 1.33 1.41
64 K 3.00 0.96 4.25 0.76 0.80
32 K 1.72 0.44 2.26 0.39 0.40
Figure 8: Comparison of maximum supported training sequence length with ZigZag-RingAttn (Li et al., 2023a; Liu et al., 2024a; Zhu, 2023; Korthikanti et al., 2023) and DeepSpeed-Ulysses (Jacobs et al., 2023). Our system supports longer training sequence lengths than Ulysses and is comparable to ZigZag-RingAttn.
Figure 9: Performance of training and inference systems on 32 and 8 H100 GPUs.
Table 3: Comparison of first token latency (left; seconds) and maximum supported inference sequence length (right) against the native HuggingFace Pipeline.
Seq. 8K 16K 32K 64K 96K 128K 256K
HF 1.81 2.47 4.20 8.50 15.36 OOM OOM
Ours 0.22 0.30 0.53 1.08 1.88 2.89 9.24
# GPUs 1 2 4 8
HF 48K 83K 96K 96K
Ours 48K 88K 138K 276K

### v4

@@ Abstract
Long-context capability is critical for multi-modal foundation models, especially for long video understanding. We introduce LongVILA, a full-stack solution for long-context visual-language models by co-designing the algorithm and system. For model training, we upgrade existing VLMs to support long video understanding by incorporating two additional stages, i.e., long context extension and long video supervised fine-tuning. However, training on long video is computationally and memory intensive. We introduce the long-context Multi-Modal Sequence Parallelism (MM-SP) system that efficiently parallelizes long video training and inference, enabling 2M context length training on 256 GPUs without any gradient checkpointing. LongVILA efficiently extends the number of video frames of VILA from 8 to 2048, improving the long video captioning score from 2.00 to 3.26 (out of 5), achieving 99.8% accuracy in 6,000-frame (more than 1 million tokens) video needle-in-a-haystack. LongVILA-7B demonstrates strong accuracy on the VideoMME benchmark, i.e., 61.8% with subtitle. Besides, MM-SP is 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than ring style sequence parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron with a hybrid context and tensor parallelism. Moreover, it seamlessly integrates with Hugging Face Transformers.
@@ 1 Introduction
While some works have enabled long-context Vision-Language Models (VLMs) (Lin et al., 2023b; Weng et al., 2024), they employ simplified approaches rather than offering a comprehensive solution. For instance, LongVA (Zhang et al., 2024a) relies on long-context LLMs and trains models on short-context data. LongVLM (Weng et al., 2024) utilizes token compression to circumvent context extension. These approaches sidestep more challenging issues, such as the development of a robust long-context multi-modal training framework and corresponding dataset design.
@@
A full-stack design is crucial for long-context Vision-Language Models (VLMs). Training large models is typically a complex, systematic endeavor that demands both data engineering (Betker et al., 2023; Ouyang et al., 2022; Zhou et al., 2024) and system-software co-design (Lepikhin et al., 2020; Chowdhery et al., 2023; Shoeybi et al., 2019; Brown et al., 2020; Dehghani et al., 2023). Unlike text-only LLMs, VLMs (e.g., LLaVA (Liu et al., 2023c)) often require distinct model architectures and flexible distributed training strategies. Additionally, long-context modeling necessitates not only long-context data to fully utilize the model’s capabilities (Fu et al., 2024c; Chen et al., 2023) but also infrastructure capable of supporting memory-intensive long-context training (Li et al., 2021; Jacobs et al., 2023; Li et al., 2023a). Therefore, a full-stack design, encompassing training pipeline and system, is indispensable for long-context VLMs.
@@
Figure 1: The LongVILA training pipeline. In Stages 1 through 3, the process starts with alignment, pre-training, and supervised fine-tuning. In Stage 4, the model undergoes mid-training context extension. Finally, in Stage 5, the model is fine-tuned for long video understanding with Multi-Modal Sequence Parallelism (MM-SP). While this example uses VILA, the process is adaptable to other VLMs. We train LongVILA models to support 1024 video frames in Stage 5 in experiments.
Figure 3: Comparison of Needle in the Long Video Haystack Experiment. The 32-frame baseline model (left) can not retrieve right needles after 32 frames. LongVA (middle) achieves 87.6% accuracy in 3,000 frames. In contrast, the LongVILA model (right), trained on 2048 frames, presents 99.8% accuracy on 6,000 frames (more than 1 million context length).
@@
In this work, we introduce LongVILA, a comprehensive solution for long-context VLMs. For training pipeline, we implement a five-stage training curriculum as Figure 1: (1) multi-modal alignment, (2) large-scale pre-training, (3) short supervised fine-tuning, (4) context extension for LLMs, and (5) long supervised fine-tuning. For system, we establish an efficient and user-friendly framework, namely Multi-Modal Sequence Parallelism (MM-SP), which supports training and inferencing memory-intensive long-context VLMs.
@@
LongVILA-7B presents strong performance on VideoMME (Fu et al., 2024a), 61.8% with subtitle. The LongVILA model, trained on 2048 frames, achieves 99.8% accuracy in the needle-in-a-haystack experiments with 6,000 frames, with a context length of more than 1 million tokens. In ablations, by increasing the number of video frames using LongVILA, the performance on VideoMME and long video captioning tasks consistently improves (Figure 2). Our MM-SP system can efficiently scale the context length up to 2 million tokens without gradient checkpointing, achieving 2.1$\times$ to 5.7$\times$ speedup compared to ring style sequence parallelism, and 1.1$\times$ to 1.4$\times$ compared to Megatron with a hybrid context parallelism and tensor parallelism.
@@ 2 Background and Related work
There are two predominant designs for VLMs: the encoder-decoder architecture (e.g.,, LLaVA (Liu et al., 2023c), PaLM-E (Driess et al., 2023)) and the decoder-only architecture (e.g.,, Fuyu (Bavishi et al., 2023), Chameleon (Team, 2024)). Encoder-Decoder VLMs connect the vision encoder to the LLM decoder through a multi-modal projector. Certain multi-modal projectors, such as spatial pooling and Q-former, significantly reduce the number of tokens per image or video frame, thereby lowering the computational burden on the LLM decoder. In contrast, decoder-only LLMs typically process raw patches as input without hierarchical token pooling, making it more challenging to reduce the token count for each image or frame. In this work, we build on VILA (Lin et al., 2023b) as our foundation. It is worth noting that enhanced variants of VILA exist, such as VILA2(Fang et al., 2024) for improved performance and X-VILA(Ye et al., 2024) for cross-modality understanding, reasoning, and generation. For our model architecture and training pipeline, we adhere to the standard VILA-1.5 version.
@@ 3 LongVILA Training Pipeline
As shown in Figure 1, in our pipeline, there are five training stages, i.e., Stage 1: multi-modal alignment, Stage 2: large-scale pre-training, Stage 3: supervised fine-tuning, Stage 4: context extension for LLM, Stage 5: long supervised fine-tuning. Stage 1, 2, and 3 follow VILA (Lin et al., 2023b), to firstly bridge the gap between LLM and vision encoder, and then pre-training on larger datasets. In Stage 1, only the multi-modal projector is trainable with others frozen. In Stage 2, we freeze the vision encoder and training LLM and the multi-modal projector. In Stage 3, we fully fine-tune the model for short data instruction following, e.g., image and short video datasets. Afterwards, we extend the context length of LLM with text-only dataset in a continued pre-training manner in Stage 4. In Stage 5, we adopt our MM-SP system (§4) to enhance the instruction following abilities by long video supervised fine-tuning. It is noted that all parameters are trainable in the final stage.
Figure 4: The proportion of question and video categories in our LongVILA__\__sft dataset. We have 15,292 videos in total. For each video, there are one sample for captioning and the other question.
@@ 3.1 Stage1&2&3: Alignment, Pre-training, and Short Supervised Fine-tuning
We first use open-sourced image and video caption datasets to train the multi-modal projector in stage (1) to conduct the multi-modal alignment. Note that, following (Lin et al., 2023b), both vision encoder and LLM decoder are frozen at this stage. After that, we conduct large-scale pre-training to learn general multi-modal capacity at scale. To improve the quality of large open-sourced datasets, we follow VILA2 (Fang et al., 2024) to relabel COYO-25M (Lin et al., 2023b; Byeon et al., 2022) with VILA-1.5-40B (Lin et al., 2023b). The supervised fine-tuning process incorporates mixed data types, including both images and videos. For short video comprehension, we utilize open-source video instruction-following datasets, e.g., YouCook2 Zhou et al. (2018) and ShareGPTVideo Zhang et al. (2024b). In experiments, our model is based on Qwen2-7B (qwe, 2024) for benchmark comparisons, e.g., VideoMME, and Llama3-8B (AI@Meta, 2024) for ablation study.
@@ 3.2 Stage4: Context Extension for LLMs
Our empirical research indicates that extending the context length of LLMs is essential prior to engaging in supervised fine-tuning with long video datasets. Following Stage 2 of our methodology, we execute a continuation of pre-training on the LLM to enhance its context length to 262,144, utilizing a total of 17B tokens. We employ a progressive training schedule, incrementally increasing the context length from 8,192 to 65,536, and ultimately to 262,144, utilizing the SlimPajama dataset (Soboleva et al., 2023) in accordance with the methodology outlined by (Fu et al., 2024d).
@@ 3.3 Stage5: Long Supervised Fine-tuning
Long video instruction following To facilitate the fine-tuning of long videos, we constructed a new, dedicated dataset for long video training, each consisting of 15,292 videos. We use the original long videos from the Shot2Story dataset (Han et al., 2023). Each video includes different questions and answers: one for generating captions and another for answering questions, enabling diverse applications in video understanding. Figure 5 illustrates the process for generating instruction-following datasets from long videos. Initially, the long video is segmented into shorter clips, each approximately 10 seconds in duration. These clips are then independently annotated with descriptive captions utilizing the VILA-1.5 model. Subsequently, an LLM is employed to generate question-and-answer pairs derived from the captions of these clips. The generated questions encompass summarization and other queries relevant to the comprehensive understanding of long video content.
@@
Once we acquired the long video dataset, applying it for supervised fine-tuning introduced new challenges, primarily due to the substantial number of frames in each sample—often ranging in the hundreds or even thousands. For instance, a single sequence from 1400 video frames can encompass around 274k tokens. Existing data-parallel training systems struggle to handle such extensive contexts. We developed the MM-SP system (Section 4) to efficiently train long-context VLMs.
@@ 4 Multi-Modal Sequence Parallelism
Figure 6: Sharding strategy and communication pattern of MM-SP. For sharding strategy, Ring SP is designed for text-only modalities, without optimization for the workload of an image encoder. Our MM-SP implements a novel sharding strategy that balances the computational load between the image encoder and the language modeling stages. For communication pattern, Ring SP (Liu et al., 2023a; Li et al., 2023a) (top) relies on P2P communication for both intra-node and inter-node settings, resulting in underutilization of intra-node bandwidth. MM-SP (bottom) adopts 2D-Attention (Fang & Zhao, 2024; Gu et al., 2024) mechanism which utilizes intra-node All-to-All (All2All) and inter-node Point-to-Point (P2P) communication to transfer keys and values (KV), enhancing the efficiency of intra-node NVLink utilization. The bandwidth is for H100.
4.1 Limitations of Existing Systems
@@ 4.2 Multi-Modal Sequence Parallelism Training Mode
Figure 7: Workflow of Multi-Modal Sequence Parallelism with a Batch Size (bs) of 4 and a Sequence Parallel Size (SP_Size) of 3. To accomodatemulti-modal inputs, we developed a customized sharding strategy that ensures balanced workload distribution and compatibility with SP communucation.
@@ 4.3 Multi-Modal Sequence Parallelism Inference Mode
The model we developed through sequence parallelism training is capable of handling long-context multi-modal downstream tasks. However, the most commonly used inference system, built on HuggingFace Transformers, typically operates on a single GPU. This lack of distributed implementation limits the maximum sequence length that can be processed during inference. The most straightforward solution within HuggingFace Transformers is to use its pipeline parallelism inference feature, which shards a single model across multiple devices on a layer-by-layer basis (Huang et al., 2019; Narayanan et al., 2019). However, this approach is inefficient, as it only activates one GPU at a time. Additionally, it struggles to support long sequence because the first device must store large input embeddings and images, creating a memory bottleneck.
To address these limitations, we implemented sequence parallelism for distributed inference in VLMs. Unlike the training mode, the inference system additionally manages tensors, such as input tokens and position encodings, that progressively change during the decoding phase (Yu et al., 2022). It detects signals from the machine with the last token to terminate the distributed process appropriately. Compared to HuggingFace’s pipeline parallelism strategy, our inference mode is more efficient, as all devices participate in computation simultaneously, accelerating the process by a factor proportional to the number of machines (Figure 8). Furthermore, it is scalable, with memory evenly distributed across devices, enabling longer sequences with additional machines.
@@ 5.1 Training and inference system
Figure 9: Performance comparison of training systems on 32 H100 GPUs. MM-SP is as scalable as ZigZag-RingAttn, and as efficient as Ulysses.
Figure 10: Performance comparison of inference systems on 8 H100 GPUs.